OpenStudy (anonymous):

Solve the differential equation Initial Value Problem: dy/dt=-y+5 , y(0)=Yo

OpenStudy (anonymous):

where Yo=\[y _{0}\]

OpenStudy (anonymous):

this is a separable differential equation\[\frac{\text{d}y}{-y+5}=\text{d}t\]integrate

OpenStudy (anonymous):

okay I did that but in the back of the textbook it says the answer is \[y=5+(y _{0}-5)e ^{-t}\]

OpenStudy (anonymous):

and I have no idea how that happened

OpenStudy (anonymous):

well we have\[\frac{-\text{d}y}{-y+5}=-\text{d}t\]\[\ln(5-y)=-t+c\]apply exp\[5-y=ae^{-t}\]

OpenStudy (anonymous):

but why did you multiply both sides by -1?

OpenStudy (anonymous):

ohhh okay nevermind

OpenStudy (anonymous):

I'm assuming a is y0?

OpenStudy (anonymous):

\[\ln(5-y)=-t+c\]apply exp\[5-y=e^ce^{-t}\]let \(a=e^c\) it does'nt matter because e^c is a constant then apply the initial value\[5-y_0=ae^0=a\]\[a=5-y_0\]

OpenStudy (adunb8):

im having same problems i guess your taking differential equation & linear algebra also.

OpenStudy (anonymous):

oh okay thanks mukushla, you helped me out a lot

OpenStudy (anonymous):

:)